# Modelling Mean-Field Games with Neural Ordinary Differential Equations

### v2

@@ Abstract
Mean-field game theory relies on approximating games that would otherwise have been intractable to model. While the games can be solved analytically via the associated system of partial derivatives, this approach is not model-free, can lead to the loss of the existence or uniqueness of solutions and may suffer from modelling bias. To reduce the dependency between the model and the game, we combine mean-field game theory with deep learning in the form of neural ordinary differential equations. The resulting model is data-driven, lightweight and can learn extensive strategic interactions that are hard to capture using mean-field theory alone. In addition, the model is based on automatic differentiation, making it more robust and objective than approaches based on finite differences. We highlight the efficiency and flexibility of our approach by solving three mean-field games that vary in their complexity, observability and the presence of noise. Using these results, we show that the model is flexible, lightweight and requires few observations to learn the distribution underlying the data.
@@ 1 Introduction
Figure 1: An overview of modelling MFGs with neural ODEs. The neural ODE combines the MFG mechanics with the neural network output to create more informed strategies.
@@
Game theory investigates strategic interactions between self-interested agents that inhabit the same environment and are affected by each other’s actions. Such analysis has applications in economics [1, 2, 3], crowd motions [4, 5] and evolutionary biology [6]. We propose neural network-based methods to solve mean-field games (MFGs), which describe games with an extremely large (or, in the limit, infinite) population of agents [7, 8]. MFG theory has been introduced by Lasry and Lions [8] and Huang et al. [7] and borrows methodologies from particle physics. Given the large number of particles in this field, it is expensive at best and intractable at worst to model all inter-particle interactions. Instead, mean-field theory replaces the particles with their mean-field, the distribution from which the theory derives its name [9]. From a mathematical perspective, this distribution is described with two coupled differential equations: the Hamilton-Jacobi equation for expressing the change in dynamics of a single particle; and the Fokker-Planck equation that describes the mean-field density summarizing the interactions between all other particles [8, 10].
When games are considered, the atomic agents of interest become more complex: in contrast to fundamental particles, players can perceive and act on their environment, which allows for strategic decision-making. To accommodate this rationality, while still being able to construct a mean-field, MFG theory relies on assumptions of symmetry and homogeneity [8].
The analysis of MFGs concerns optimality conditions that arise from the joint minimization of the agents’ costs. The optimal control leading to such a minimum is found by solving the partial or stochastic differential equations (DEs) accompanying the MFG [10]. These systems can be solved numerically by discretizing the differential equations using finite differences [11, 12]. However, solving a system of differential equations involving high-dimensional agent states is challenged by the curse of dimensionality [13, 14]. In addition, the discretization introduces an error that is subject to the chosen method and may ultimately lead to the loss of the existence or uniqueness of the solutions [15].
@@
Instead, the solutions to the mean-field game can be approximated using artificial neural networks [16], in the form of recurrent architectures [17, 18], deep policy iteration [19], deep Galerkin methods [20] or actor-critic-based approaches [21, 15, 22]. Still, all approaches mentioned above are DE-based and require full knowledge of the game of interest. This dependency limits the models’ applications and may introduce a modelling bias. Reinforcement learning (RL)-based solutions surmount this requirement by building upon the Fictitious Play algorithm instead of following a numerically exact approach [23, 24, 25, 26, 27]. Because the Fictitious Play algorithm is independent of the modelled system, the RL-based algorithms are commonly referred to as “model-free" [13].
@@
To retain the numerical exactness of the DE-based solutions without requiring full knowledge of the game of interest, we propose to analyze MFGs using neural ordinary differential equations (neural ODEs; Chen et al., 2018). To do so, we complement the existing system of DEs with a neural network (Figure 1). The resulting model is data-driven and may therefore alleviate modelling bias associated with the predefined MFG dynamics. Furthermore, the model’s ability to recover structure from data allows it to reason beyond the predefined DEs. Lastly, the neural ODEs are solved using automatic differentiation, which makes them more robust than the DE-based solutions based on discretization [29]. This property is, in combination with the model’s data-driven nature, expected to prevent the loss of the existence of solutions that affect the original PDE-based approaches.
@@
Contributions   To the best of our knowledge, we are the first to combine mean-field mechanics with deep learning in this way. Therefore, the contributions of this paper are threefold: (1) We provide the theoretical framework for combining MFG theory with neural ODEs, (2) we reformulate examples from classical game theory to their MFG counterparts, taking the continuum limit on the number of agents by leveraging the structure of the neural ODEs, and (3) we test the neural ODEs on the games presented in Table 1 to give a proof of concept of the model’s ability to construct new (data-driven) Nash equilibria by empirically highlighting emerging strategies.
Table 1: An overview of the mean-field games included in the Experiments.
Game Observability Neural network input Neural network output Meeting arrival times Full Observed arrival times Intended arrival times El Farol Bar problem Full Observed attendance rates Probability of going to the bar Liars dice Partial Dice outcomes, dice odds and bluff strategy Dice odds and bluff strategy
@@ 2 Background
We combine the concepts of mean-field games with neural ODEs. We use bold characters to denote vectors.
2.1 Mean-Field Games
Depending on the context, the MFGs are commonly described using partial differential equations (PDEs) or concepts from reinforcement learning. This section considers the PDE perspective, making it easier to unite the MFGs with the neural ODEs discussed later.
To introduce the system of PDEs describing the game, consider a system of $N$ agents. Each agent $i\in[1,\dots,N]$ has a state at time $t\in[t_{0},T]$, denoted with $\bm{x}_{t}^{i}\in\mathbb{R}^{d}$. The state of agent $i$ evolves over time according to their control, or game-playing strategy, $\bm{\alpha}_{t}^{i}$. The dynamics of a single agent can be described according to the Itô process presented in Equation 1, combining their drift function $b$ with a diffusion driven by a Brownian motion. The strategies of all other agents are accounted for with the inclusion of $m$, a probability measure representing the mean-field [25]. $\nu$ is a nonnegative parameter.
$\frac{d\bm{x}}{dt}=b\left(\bm{x}_{t}^{i},\bm{\alpha}_{t}^{i},m(t)\right)+\sqrt{2\nu}B_{t}$ (1)
The agents’ state evolves by the repeated application of their drift and diffusion, such that $\bm{x}_{t+1}=\bm{x}_{t}+b\left(\bm{x}_{t}^{i},\bm{\alpha}_{t}^{i},m(t)\right)+\sqrt{2\nu}B_{t}$. Having the freedom to optimize their control in $\mathbb{R}^{k}$, the agents minimize the cost function $J$.

### v3

# Neural Mean-Field Games: Extending Mean-Field Game Theory with Neural Stochastic Differential Equations

@@ Abstract
Mean-field game theory relies on approximating games that are intractible to model due to a very large to infinite population of players. While these kinds of games can be solved analytically via the associated system of partial derivatives, this approach is not model-free, can lead to the loss of the existence or uniqueness of solutions, and may suffer from modelling bias. To reduce the dependency between the model and the game, we introduce neural mean-field games: a combination of mean-field game theory and deep learning in the form of neural stochastic differential equations. The resulting model is data-driven, lightweight, and can learn extensive strategic interactions that are hard to capture using mean-field theory alone. In addition, the model is based on automatic differentiation, making it more robust and objective than approaches based on finite differences. We highlight the efficiency and flexibility of our approach by solving two mean-field games that vary in their complexity, observability, and the presence of noise. Lastly, we illustrate the model’s robustness by simulating viral dynamics based on real-world data. Here, we demonstrate that the model’s ability to learn from real-world data helps to accurately model the evolution of an epidemic outbreak. Using these results, we show that the model is flexible, generalizable, and requires few observations to learn the distribution underlying the data.
Keywords
Mean-field game theory, Neural stochastic differential equations, Nash Equilibria, Compartmental models
@@ 1 Introduction
Game theory investigates strategic interactions between self-interested agents that inhabit the same environment and are affected by each other’s actions. Such analysis has applications in economics [35, 32, 5], crowd motions [4, 13] and evolutionary biology [7]. Some of these applications may consider an extremely large, or in the limit, infinite population. As it is intractable to model all interactions between such a large number of agents, these types of games are approximated using mean-field game (MFG) theory [51, 40]. MFG theory replaces the agents with their mean field, effectively summarizing their interactions by a distribution [37]. From a mathematical perspective, this distribution is described with two differential equations that are coupled through the agents’ cost function: the first, a Hamilton-Jacobi-Bellman equation, guides the optimization strategy of the average agent, whereas the second, a Fokker-Planck equation, describes the temporal evolution of the population’s overall dynamics.
@@
Even though each agent influences the population’s overall dynamics by minimizing their own expected costs, the infinitesimal agents are too small to independently influence the population. Instead, all agents are expected to jointly converge towards an optimal control that minimizes their individual costs. This state, which can be considered the game’s Nash Equilibrium, can be found by numerically solving the partial or stochastic differential equations that accompany the game [21]. This can, for example, be done by numerically solving the differential equations using finite differences [18, 1]. Even so, solving a system of differential equations involving high-dimensional agent states is challenged by the curse of dimensionality [52, 16]. In addition, the discretization introduces an error that is subject to the chosen method and may ultimately lead to the loss of the existence or uniqueness of the solutions [70]. Instead, the solutions to the mean-field game can be approximated using artificial neural networks [54], in the form of recurrent architectures [24, 25], deep policy iteration [11], deep Galerkin methods [10], or actor-critic-based approaches [33, 70, 69]. Still, all approaches mentioned above require full knowledge of the game of interest. This dependency limits the models’ applications and may introduce a modelling bias. Reinforcement learning (RL)-based solutions surmount this requirement by building upon the Fictitious Play algorithm instead of following a numerically exact approach [66, 53, 60, 8, 58]. Because the Fictitious Play algorithm is independent of the modelled system, the RL-based algorithms are commonly referred to as “model-free” [52].
@@
To retain the numerical exactness of the differential equation-based solutions without requiring full knowledge of the game of interest, we propose to analyze MFGs using neural stochastic differential equations (SDEs; [68, 55]). Neural SDEs form the stochastic extension of neural ordinary differential equations (ODEs), a family of neural networks introduced by Chen et al. [28]. To model the MFG dynamics as a neural SDE, we complement the existing system of coupled differential equations with a neural network. The resulting model, shown in Figure 1, is data-driven and may therefore alleviate the modelling bias associated with the predefined MFG dynamics. Furthermore, the model’s ability to recover structure from data allows it to reason beyond the predefined differential equations. Lastly, the neural SDEs are solved using automatic differentiation, which makes them more robust than approaches driven by classical numerical methods such as finite difference and finite elements [65]. This property, in combination with the model’s data-driven nature, is expected to prevent the loss of the existence of solutions that affect the original differential equation-based approaches. We demonstrate the flexibility of the neural MFGs by modelling the dynamics of COVID-19 infections. Dynamical systems for epidemics are highly nonlinear and depend on various features, including population parameters, restrictive measures, and viral characteristics [56]. Because it can be challenging to quantify the effects of such features, their inclusion is often beyond the scope of classical mathematical models [56]. To provide insight into the complex mechanics behind epidemic systems, we extend the classical models by learning viral mechanics from a large range of epidemic characteristics.
Figure 1: An overview of modelling MFGs with neural SDEs. The resulting neural mean-field game combines the MFG mechanics with the neural network output to create more informed strategies.
@@
Contributions The contributions of this paper are fourfold: (1) We introduce neural MFGs, providing the theoretical framework for combining MFG theory with neural SDEs, (2) we reformulate examples from classical game theory to their MFG counterparts, taking the continuum limit on the number of agents by leveraging the structure of the neural SDEs, and (3) we test the neural SDEs on the two games with increasing complexity to give a proof of concept of the model’s ability to construct new, data-driven Nash equilibria by empirically highlighting emerging strategies. Lastly, (4) we apply the neural SDEs on a real-world example, by augmenting a compartmental model of the COVID-19 outbreak in Japan and incorporating auxiliary observations.
Data and Code Availability The work uses data obtained from The Japanese Ministry of Health, Labour and Welfare [67] and Mathieu et al. [57]. The code associated with the model implementation and analyses is available on GitHub: https://github.com/KachmanLab/NeuralMFG.
@@ 2 Background
In this section, we review MFG theory and the neural ODEs and SDEs introduced in [55, 28]. MFG theory, which has largely been shaped by a series of contributions by Lasry and Lions [49, 50, 51], proposes an approximation of the Nash equilibria for stochastic differential games with a large number of players [23]. As mentioned before, MFGs can be described using both partial differential equations (PDEs) and concepts from RL. This section considers the PDE perspective, which allows for an analytical and tractable formulation of MFGs. In addition, treating the formulation based on differential equations makes it easier to unite MFGs with the neural SDEs discussed later.
2.1 Modelling Games with an Infinite Number of Players
MFGs model the dynamics of the infinitesimal players and the full population by jointly solving the Hamilton-Jacobi-Bellman (HJB) equation and the Fokker-Planck (FP) equation. To illustrate this coupling, we consider a stochastic differential game with $N$ players. Each player $i\in[1,\dots,N]$ has a state at time $t\in[t_{0},T]$, denoted with $x_{t}^{i}\in\mathbb{R}^{d}$. The state of agent $i$ evolves over time according to $\alpha_{t}^{i}$: the player’s control, or game-playing strategy, at time $t$. The control is a function that takes values within the set of admissible controls $\mathcal{A}$. We assume that the dynamics of a single player are given by a (general) Itô process [72, p. 21] of the form
$dx^{i}_{t}=b\left(t,x_{t}^{i},\alpha_{t}^{i},m(t)\right)\,dt+\sigma(t,x_{t}^{i},\alpha_{t}^{i},m(t))\,dB_{t}$ (1)
where the deterministic and stochastic terms $b$ and $\sigma$ respectively describe the player’s drift and diffusion, and $B$ is an Wiener process (Brownian motion). Naturally, the individual players may base their strategy on other players’ actions. This dependency is accounted for with the inclusion of $m(t)=\frac{1}{N}\sum^{N}_{i=1}\delta_{x_{t}^{i}}$, the empirical distribution representing the mean-field of all other players’ states $x_{t}=(x_{t}^{1},\dots,x_{t}^{N})$ [21, 23]. In this formulation, $\delta_{x}$ denotes the Dirac measure at $x$.
2.2 Mean-Field Game Principles
An MFG can be seen as an optimization problem, where each player chooses a strategy that minimizes their expected cost over the period $t_{0},T]$. This agent-specific cost, $J^{i}$, is given by the quantity
$J^{i}=\mathbb{E}\left[\int_{t_{0}}^{T}\frac{1}{2}L(x_{s}^{i},m(s),\alpha_{s}^{i})ds+G(x_{T}^{i},m(T))\right].$ (2)
Following the fundamentals of dynamic programming, $J$ combines the evolution of the player’s cost $L$ through $s$ with a terminal cost $G$ at the state $T$ [21]. Just like the agents’ dynamics, the cost function links the agents’ individual states $x$ to the mean-field $m$. Assuming that agent $i$ is a representative agent in the system, it is implied that all agents aim to minimize the cost function. Therefore, MFGs are commonly analyzed with the generalized notion of the Nash equilibrium, which is reached for a family of controls ($\bar{\alpha}\in\mathcal{A})$ when
$\forall i,j\in[1,\dots,N],\;J^{i}\left(\bar{\alpha}^{i},\bar{\alpha}^{j}_{j\neq i}\right)\leq J^{i}\left(\alpha^{i},\bar{\alpha}^{j}_{j\neq i}\right)$ (3)
holds for all non-Nash strategies $\alpha^{i}$ [61]. That is, the Nash equilibrium is a fixed point in the space of the flow in $m$ [26], which implies that no agent benefits from changing their strategy if a Nash equilibrium is reached. Computing the Nash equilibria for many agents involves a large set of controls and can be a lot of work. Fortunately, MFG theory simplifies this problem with the assumptions of symmetry and homogeneity [37], facilitating the calculation of Nash equilibria when $N\to\infty$. In short, these assumptions imply that all players $i$ are statistically identical and can be described with the same drift $b$ and diffusion $\sigma$ [23]. We refer the reader to Cardaliaguet [21] for a detailed account of this simplification.
To facilitate the approximation of the Nash equilibria when $N\to\infty$, Lasry and Lions [51] formalize MFGs as a system of PDEs, as shown in Equation 4. The system consists of (a) a HJB equation providing the necessary and sufficient conditions for the optimality of $\alpha$ w.r.t. $J$ and (b) a FP equation describing the evolution of the mean-field according to the dynamics presented in Equation 1.
$\begin{cases}\text{(a)}\quad-\partial_{t}u-\nu\Delta u+H(x_{t}^{i},m(t),Du)=0&\text{in $\mathbb{R}^{d}\times(0,T)$}\\ \text{(b)}\quad\partial_{t}m(t)-\nu\Delta m(t)-\text{div}\left(D_{p}H\left(x_{t}^{i},m(t),Du\right)m(t)\right)=0&\text{in $\mathbb{R}^{d}\times(0,T)$}\\ \text{(c)}\quad m_{0}=m(0),\;u(x,T)=G(x_{T}^{i},m(T))\end{cases}$ (4)
The coupling of the HJB and FP equations is facilitated via the unknowns $u$ and $m$. In particular, the first equation goes backwards in time and describes the value function of a single agent, uu.111uu can be considered the supremum of the optimization problem described by the functional $J$, that is, $u(x,t)=\underset{\alpha\in\mathcal{A}}{\sup}\;J(\alpha)$ [2, 22]. The second goes forward in time and describes the change in the mean-field density $m(t)$. The equation is parameterized with the nonlinear Hamiltonian $H$, which is the Fenchel conjugate to $L$ w.r.t. the mean-field probability $p$. In addition, $D$ denotes a differential operator and $\nu$ is a nonnegative parameter. Both equations are subject to two structure conditions [51]: Firstly, $H(x,m,p)$ is convex with respect to pp. [^2] This is required to satisfy the optimality condition provided by the HJB equation [51]. Secondly, $m$ is a probability measure with the initial state $m(0)$ (Equation 4c). In addition to the structure conditions, the terminal condition $u(x,T)=G(x_{T}^{i},m(T))$ relates Equation 4 back to the cost presented in Equation 2 [51, 21].
2.3 Neural Differential Equations
Neural ODEs, and neural SDEs in their extension, are a neural network-based method introduced by Chen et al. [28]. To introduce the method, we consider a simple ODE $h$ describing the drift $b$ of $x$ over time $t$.
h(t,x(t))=b(t,x)dt.h(t,x(t))=b(t,x)\,dt. (5)
To learn the development of the dynamics of $x$, we can solve the initial value problem (IVP) consisting of the initial state $x(t_{0})=x_{0}$ and the ODE describing their dynamics $\nicefrac{{dx}}{{dt}}=h(t,x(t))$. The solution to this problem is commonly calculated using discretization, transforming the continuous function $h$ into a counterpart that consists of small steps [6]. In this case, $x(t)$ is updated step-by-step by adding its identity to $h(t,x(t))$, the output of the ODE. This process is illustrated in Equation 6, where $\Delta t$ denotes the discretized stepsize.
$x(t+\Delta t)=x(t)+h(t,x(t))$ (6)
Chen et al. [28] highlight the similarity between the discretization of continuous functions and the residual motif $y=F(x)+x$ from ResNet architectures [39]. They compare the addition of the identity $x(t)$ to the residual skip connections, relating the discretization steps to the repeated application of the residual blocks $F$. This analogy lays the foundation for neural ODEs: a class of differential equations expressed as a neural network. Thus, instead of discretizing a concrete mathematical function $h$, the neural ODE repeatedly applies the data-driven output of a neural network $f_{\theta}$. The discretized step size, $\Delta t$ is optimized by the differential equation solver and may vary throughout the solving process. As a consequence, a neural ODE can be seen as a neural network with a flexible number of layers. The output of the solver can be regarded as a latent trajectory over time $t\in[t_{0},T]$, where each output $\bm{z}_{t_{0}},\dots,\bm{z}_{T}$ represents the latent state [28]. Because the neural ODEs are trained using reverse automatic differentiation and the adjoint sensitivity method [62], the user has explicit control of the model’s numerical error [28].
2.3.1 Modelling Real-World Stochastic Processes
Instead of merely modelling a deterministic function, neural SDEs extend the neural ODEs to deal with the presence of stochasticity and randomness underlying the process [68, 41, 59]. In a neural SDE (Equation 7), both the drift $\mu_{\theta}$ and diffusion $\sigma_{\theta}$ are expressed as a neural network. Thus, instead of learning a single path, a neural SDE learns the distribution over path-valued random variables. The existence and uniqueness of a strong solution for $y$ is guaranteed as long as (1) $\mu_{\theta}$ and $\sigma_{\theta}$ are Lipschitz, which typically holds for neural networks, and (2) $\mathbb{E}[y(0)^{2}]<\infty$ [44]. In Equation 7, which shows the neural SDE dynamics, $\theta$ is used to refer to the neural network parameters.
$dy(t)=\mu_{\theta}(t,y(t))\,dt+\sigma_{\theta}(t,y(t))\,dB_{t}$ (7)
Because both the neural ODEs and SDEs are solely specified in terms of neural networks, the methods are data-driven. All the while, dynamical systems may be described using domain knowledge based on scientific models or physical laws. Rackauckas et al. [63] unify the above with the introduction of universal differential equations (UDEs), combining the output of $f$ with that of a concrete ODE $h$: $\nicefrac{{dx}}{{dt}}=h(t,x(t))+f(x)$. In this case, $f$ can be trained to yield insights that go beyond the predefined ODE $h$. A stochastic UDE variant has already been used to solve an HJB equation, highlighting the model’s adaptivity, efficiency, and flexibility towards stiffness [63].
3 Neural Mean-Field Games
We introduce neural Mean-field games, a new perspective on UDEs that combines existing MFG specifications with neural SDEs. To do so, we include the neural SDE at the agent level. This way, the neural network output may provide additional insights drawn from the mean-field, past game states, or observational data. Because these insights directly inform the agents’ value function $u$, the agents are expected to make better-informed decisions. In addition, this approach leverages the setup of the neural SDE, which allows for taking the number of agents or turns to the continuum limit.
3.1 Data-Driven Dynamics
The addition of neural networks to the MFG dynamics has implications for the formulation. In particular, we update the description of the single-player dynamics (Equation 1) by adding the neural drift from Equation 7 and modelling the diffusion in a fully data-driven manner. This results in the general formulation of the neural MFG dynamics, presented in equation 8.
$dx^{i}_{t}=\left[b\left(t,x_{t}^{i},\alpha_{t}^{i},m(t)\right)+\mu_{\theta}(t,x_{t}^{i},a_{t}^{i},m(t))\right]\,dt+\sigma_{\theta}(t,x_{t}^{i},\alpha_{t}^{i},m(t))\,dB_{t}$ (8)
We make the neural MFG data-driven by incorporating a loss term that minimizes the discrepancy between the players’ states and observations. In this case, we extend the cost function $J^{i}$ with a loss term comparing the predicted distribution at time to the observed label. This loss term can be included in the running costs $L$, or in the terminal costs $G$, depending on the MFG of interest. Thus, in contrast to a single MFG, a neural MFG minimizes the sum of two losses: one originating from the game description and one minimizing the discrepancy between the model predictions and observed data. We will provide an interpretation of the loss in the next section. Because of the symmetry assumption and the joint minimization of the players’ loss, the game’s Nash Equilibria now depend on $b$, $\mu_{\theta}$, and $\sigma_{\theta}$ instead of only on $b$ and $\sigma$. As a result, the Nash Equilibria become data-driven.
3.1.1 A Note on the Existence and Uniqueness of
Solutions
Naturally, the inclusion of the neural networks has implications for the existence and uniqueness of the solutions to the MFG. However, we show that the formalism behind the existence and uniqueness of MFG solutions can be extended to the neural MFG case. We consider two types of MFGs: games with a continuous state space, as introduced by [51] and Huang et al. [40], as well as games with a discrete state space, as presented by Doncel et al. [29]. The proof of existence and uniqueness of continuous MFG solutions is based on Picard’s existence theorem [20, Theorem 110C, and Appendix A] and is well-established in the literature, e.g. [42, 3, 23, 21]. Similarly, the existence and uniqueness of solutions for continuous-time, finite-state MFGs have also been thoroughly considered [36, 15, 27]. For the discrete MFG considered in Section 4.3, the proof is provided by Doncel et al. [30] and is based on Carathéodory’s Existence Theorem, which can be applied to discontinuous differential equations [17, Appendix A]). As both proofs rely on the Lipschitz continuity of the functions $F$ and $G$, we can extend them to the data-driven case by assuming $\mu_{\theta}$ and $\sigma_{\theta}$ to be Lipschitz continuous as well. [^3] This is a fair assumption to make as neural networks are generally formed by a composition of Lipschitz functions [44]. As long as the outputs of the neural networks are added to the MFG dynamics, the resulting neural MFG consists of a sum of Lipschitz functions, which is again Lipschitz continuous.
3.2 Implementation
Even though it might be tempting to supply each agent with their own network to increase the diversity of strategies, this approach is not in line with the assumption of symmetry and drastically reduces the scalability of the method. Thus, we use a single neural network per MFG. More specifically, we use a multilayer perceptron with 3 to 8 hidden layers of 8 to 32 neurons. Even though the network seems very small, we stress that it’s applied repeatedly during the differential equation solve, which increases its capacity. The network uses LipSwish activations to impose the Lipschitz condition [46] and is optimized with Adabelief [71]. The learning rate is $5\times 10^{-4}$. The experimental framework has been implemented in Python. We use Optax, Diffrax, and Equinox to connect the methods to the JAX ecosystem [44, 45, 14]. All experiments have been performed on an Apple M1 Pro CPU with 16 GB of memory.
4 Experiments
We demonstrate the capacity and flexibility of the neural MFGs by solving two example games. We start with the meeting arrival times game [37], a toy example considering a distribution of (unbounded) arrival times. With the El Farol Bar problem [9], we discuss a more complex game that requires the estimation of probabilities. Next, we use the framework to simulate a real-world problem by considering a compartmental model for simulating the dynamics of COVID-19 in Japan. For each game, we extend the modelled drift with a neural network and describe how that influences the cost function associated with the MFG.
4.1 Meeting Arrival Times
The meeting arrival times game, introduced by Guéant et al. [37, p. 9], illustrates a simple MFG application. The game considers a meeting that is scheduled at time $s$. Despite this planning, the true starting time of the meeting $\tilde{s}$ depends on the fraction of agents that have arrived. We assume that $\tilde{s}\geq s$, meaning that the meeting starts later if the quorum has not been reached at time $s$. Each agent $i$ has an intended arrival time $\tau^{i}$ and actual arrival time $\tilde{\tau}^{i}$. The agents’ control $\tau$, but $\tilde{\tau}^{i}$ is subject to the uncertainty $\tilde{\epsilon}^{i}$ such that $\tilde{\tau}^{i}=\tau^{i}+\tilde{\epsilon}^{i}$. In our implementation of the game, $\tilde{\epsilon}^{i}$ is drawn from the Gaussian distribution $\mathcal{N}\left(0,1\right)$.
To build upon the formalism introduced in Section 2.1, consider that the agents optimize their intended arrival time by minimizing their cost $J^{i}$. We consider the variant of the game that only has a terminal cost $G$, which is the sum of the components $g_{[1-3]}$ presented below. The notation $\cdot$$_{+}$ is shorthand for $\text{max}(\cdot,0)$.
  1. $g_{1}$
    $g_{1}(s,\tilde{\tau}^{i})=\left[\tilde{\tau}^{i}-s,0\right]_{+}$
  2. $g_{2}$
    $g_{2}(\tilde{s},\tilde{\tau}^{i})=[\tilde{\tau}^{i}-\tilde{s},0]_{+}$
  3. $g_{3}$
    $g_{3}(\tilde{s},\tilde{\tau}^{i})=[\tilde{s}-\tilde{\tau}^{i},0]_{+}$
Figure 2: The distribution of meeting arrival times for the standard version of the meeting arrival times game. The randomly initialized distribution of arrival times quickly converges to a narrow distribution centred at $\tilde{s}$. The value of $\tilde{s}$ is subject to the Brownian noise of the SDE.
The behaviour of all other agents is summarized in the actual starting time $\tilde{s}$, which is the mean-field of this game that considers all agents’ actions in a cumulative distribution of arrival times. To acknowledge the coupling between $\tilde{s}$ and the individual agents, consider that each agent bases their intended arrival $\tau^{i}$ time on $\tilde{s}$ and $\tilde{s}$ is influenced by the cumulative distribution of all actual arrival times $\tilde{\tau}^{i}$. In addition, Guéant et al. [37] prove that $\tilde{s}$ exists, being a unique fixed point representing the game’s equilibrium.
We model the “standard” version of the game with an SDE that differentiates over the turns $t$ (see Equation 1). Even though the game lacks a terminal condition, we treat it as a finite game by setting $t\in[1,15]$. Figure 2 shows an example outcome, where the distribution of arrival times quickly narrows down. This observation is in accord with the proof provided by Guéant et al.. The exact value of $\tilde{s}$ diffuses between turns due to the Brownian motion in the SDE. While the original game description does not include Brownian motion, it is included in this example to demonstrate the modelling capacity of the neural MFG.
4.1.1 Learning the Intended Arrival Times
To apply the neural SDE, we create a data-driven challenge and aim to find the meeting time $s$ that aligns with a collection of observed arrival times. To this end, we convert the SDE into a neural SDE by adding the output of a feedforward neural network $f$. $f$ is trained to minimize the difference between the modelled arrival times at turn 15 and the “true” observed arrival times, $\tilde{\tau}^{*}$. We train on 10 games. For each game, $\tilde{\tau}^{*}$ is generated by drawing $M$ samples from $N$ different Gaussians $\mathcal{N}\left(\mu_{i},\sigma_{i}\right)$ for $i\in[1,2,\dots,N]$. Each $\mu_{i}$ is sampled from $\mathcal{N}(15,0.5)$, while $\sigma_{i}\sim\Gamma(2,0.2)$. As a result, $\tilde{\tau}^{*}$ consists of $N\times M$ observations that form a noisy reflection of the distribution of intended arrival times around 15:00. The neural SDE models the interplay between the game rules and the observed data by minimizing the neural network loss, which can be seen as the evolving cost function $L$, alongside the agents’ terminal cost $G$. The results are shown in Figure 3. The initial random distribution is centred around 12:00. Despite the small network size and the Brownian motion, the neural MFG converges around $s=\text{15:00}$ in turn 15.
4.2 The El Farol Bar Problem
Even though the meeting time arrival game provides a proof of concept, the neural SDE does not yield additional insights into the game dynamics. Therefore, we consider an implementation of the El Farol Bar problem, a more complex example introduced by Arthur [9]. In this problem, a fixed population of agents can go to a bar. Each agent $i$ goes to the bar with a probability $p^{i}$, and all agents need to decide simultaneously whether they’re going. The agents’ utility depends on the bar’s crowdedness. If the bar is too crowded, the agents who went there regret their decision and have less fun than the ones who stayed at home.
The cost function $J$ associated with the bar problem is comparable to that of the meeting arrival times game. However, the new cost function is selective; not all components apply to all agents. This selectivity is presumed to lead to an interplay between the MFG dynamics and the neural network. The cost is parameterized by the agent’s probability of going to the bar ($p^{i}$), the fraction of agents that went to the bar ($a$), and the crowding threshold ($c$). $J^{i}$ consists a terminal cost $G$ at state $T=15$, which is the sum of $g_{[1-3]}$ described below.
  1. $g_{1}$
    Missed a good evening The cost associated with staying at home, while the bar was not too crowded. This cost only applies to agents who did not go to the bar.